# Destructive and constructive RIS beamforming in an ISAC-multi-user MIMO network

### v1

@@ Abstract
Integrated sensing and communication (ISAC) has already established itself as a promising solution to the spectrum scarcity problem, even more so when paired with a reconfigurable intelligent surface (RIS) as RISs can shape the propagation environment by adjusting their phase-shift coefficients. Albeit the potential performance gain, a RIS also poses a security threat to the system: in this paper, we explore both sides of the RIS presence in a multi-user MIMO (multiple-input multiple-output) ISAC network. We first develop an alternating optimization algorithm, obtaining the active and passive beamforming vectors maximizing the sensing signal-to-noise ratio (SNR) under minimum signal-to-interference-plus-noise ratio (SINR) constraints for the communication users and finite power budget. We also investigate the destructive potential of RIS by devising a RIS phase-shift optimization algorithm that minimizes sensing SNR while preserving the same minimum communication SINR previously guaranteed by the system. We further investigate the impact of the RIS’s individual element failures on the system performances. The simulation results show that the RIS performance-boosting potential is as good as its destructive one and that both of our optimization strategies show some resilience towards the investigated impairments.
@@ I Introduction
Reconfigurable intelligent surfaces have captured significant interest from both academia and industry for their ability to smartly control challenging radio propagation environments. By using low-cost passive reflecting elements that induce a controllable phase-shift to incoming waveforms, RISs can effectively steer reflected signals towards desired locations while reducing interference at undesired points [1, 2]. Moreover, RISs have recently emerged as a promising technology in integrated sensing and communications (ISAC) applications, increasing the available spatial degrees of freedom. In [3], the sum-rate of the communication user equipments is maximized under the worst-case sensing signal-to-noise ratio (SNR) constraint. On the other hand, [4] defines a sensing signal-to-interference-plus-noise ratio (SINR) by taking into account the presence of interfering objects along the sensing path: this sensing SINR is then maximized in a weighted fashion alongside minimizing the communication multi-user interference.
Despite their cost-effective deployment and potential to enhance wireless communication and sensing performance, RISs present critical physical layer security (PLS) challenges. In particular, RISs could potentially be exploited as malicious entities that instead compromise communication and sensing performances. [5] gives an overview of the possible ways in which an illegally deployed RIS can hinder system performance, characterizing signal leakage from the legitimate system to the said RIS and the potential damage of interference attacks launched by the former. The addition of artificial noise onto the legitimate system’s transmitted waveform proves to be an effective strategy in increasing the system’s PLS, even though the impact of an illegally deployed RIS is still non-negligible.
@@
On the other hand, [6] assesses the impact of an active RIS, i.e., a RIS capable of amplifying the impinging signal, onto an Internet of Things network where the RIS aims at minimizing a single device’s SNR. Interestingly, the authors analyze both the cases of known and unknown malicious RIS identity.
A significant challenge is that the attacks are difficult to detect because they are not carried out by generating interfering signals, as in traditional jamming. Instead, they involve configuring the RIS to degrade the system’s performance. In [7], the impact of these silent attacks is demonstrated, further showing how destructive beamforming is susceptible to channel state information (CSI) uncertainties. An important albeit often overlooked practical consideration about RIS is that individual elements, also known as pixels, may fail: [8] gives an overview of the different types of errors that may arise in a practical RIS implementation and characterizes their impact on the RIS radiation pattern. In [9], the authors propose a failure model to specify the amplitude and phase-shift of faulty elements and they propose a diagnostic method.
In this paper, we consider a multi-user (MU)- multiple-input multiple-output (MIMO) network equipped with an ISAC base station (BS), serving multiple UEs while sensing a target. We first characterize the potential positive impact of a RIS deployment by devising an optimization problem computing active and passive beamforming vectors that maximize the sensing SNR under UEs’ SINR constraints and a finite power budget. Inspired by [10], this problem is solved through alternating optimization. We then investigate the RIS destructive potential when, unbeknownst to the BS, the sensing SNR is minimized while the same minimum SINR is still guaranteed to the UEs, as to make this attack harder to detect. We further define a pixel fault model and asses its impact on both optimization algorithms. Numerical simulations show that the RIS is equally capable of boosting and degrading the system performances and the former shows some resilience towards RIS pixel impairments.
@@ II System Model
We consider the RIS-assisted MU-MIMO ISAC network shown in Fig. 1, where a BS, equipped with $M$ transmit and $M$ receiving antennas and possessing digital beamforming capabilities, serves $K$ single-antenna UEs in the downlink while sensing the presence of a target. The system is equipped with a RIS made of $N$ reflective elements, whose phase-shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1}\dots\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$. The static path between the BS and UE $k$ is denoted by $\mathbf{h}_{s,k}\in\mathbb{C}^{M}$ whereas the static path between the BS and target is denoted by $\mathbf{h}_{s,{\rm T}}\in\mathbb{C}^{M}$. The reflected paths are defined in a similar fashion, with $\mathbf{h}_{r,k}\in\mathbb{C}^{N}$ connecting the RIS and UE $k$ while the target is connected to the RIS via $\mathbf{g}\in\mathbb{C}^{N}$. Lastly, the BS is connected to the RIS through the channel $\mathbf{H}_{t}\in\mathbb{C}^{N\times M}$. The end-to-end channel between the BS and UE $k$ is defined as
$\mathbf{h}_{k}=\mathbf{h}_{s,k}+\mathbf{H}_{t}^{\mathsf{H}}\underbrace{\text{diag}(\boldsymbol{\theta})}_{\triangleq\boldsymbol{\Theta}}\mathbf{h}_{r,k}.$ (1)
@@
where $\mathbf{f}_{l}\in\mathbb{C}^{M}$ is the precoding vector associated with the $l$-th stream. We assume that the information meant for UE $k$ is encoded into stream $k$, while an additional stream $K+1$ is used for sensing purposes. Hence, the total number of streams is $L=K+1$. On the other hand, $t_{l}$ is the complex-valued unit-power information or sensing symbol and $w_{k}$ is the complex Gaussian receiver noise with variance $\sigma_{k}^{2}$. As for the sensing performance, the two-way sensing channel is
$\mathbf{H}_{\rm T}=c_{r}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\underbrace{\mathbf{g}\mathbf{g}^{\mathsf{H}}}_{\triangleq\mathbf{T}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}+c_{s}\underbrace{\mathbf{h}_{s,{\rm T}}\mathbf{h}_{s,{\rm T}}^{\mathsf{H}}}_{\triangleq\mathbf{H}_{s,{\rm T}}}.$ (3)
The radar cross section (RCS) associated with the two sensing paths, i.e., the reflected and the static path, are modelled according to the Swerling-I model, that is $c_{r}\sim\mathcal{CN}(0,\delta_{r}^{2})$ and $c_{s}\sim\mathcal{CN}(0,\delta_{s}^{2})$. The monostatic sensing observation can be modeled as
@@
where $\mathbf{F}=[\mathbf{f}_{1}\dots\mathbf{f}_{L}]$ , $\mathbf{t}=[t_{1}\dots t_{L}]^{\top}$, and the entries of $\mathbf{w}_{\rm T}$ are independent and distributed as $\mathcal{CN}(0,\sigma_{\rm T}^{2})$. As for the RIS’s pixel malfunctioning, we adopt the clustered-biased model defined in [8]: We assume that a set $\mathcal{Q}$ of neighboring pixels assumes a phase-shift that is at a fixed distance $\kappa$ from the required one. This hardware failure effect can be taken into account by defining a failure mask $\mathbf{m}=[m_{1}\dots m_{N}]^{\top}$ as [11]
$m_{n}=\begin{cases}e^{j\kappa}\quad\text{if}\leavevmode\nobreak\ n\in\mathcal{Q}\\ 1\quad\text{otherwise}\end{cases}.$ (5)
@@ III Legitimate sensing SNR Maximization
The legitimate reason for deploying the RIS is to improve the system performance. Hence, the BS and RIS aim at jointly maximizing the sensing SNR while guaranteeing a minimum SINR to the communication UEs. The sensing SNR can be defined as
$\rho=\frac{1}{\sigma_{\rm T}^{2}}\sum_{l=1}^{L}\underbrace{\mathbb{E}\Big{[}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{\rm T}^{\mathsf{H}}\mathbf{H}_{\rm T}\mathbf{f}_{l}\Big{]}}_{\rho_{l}}.$ (6)
Similarly, the SINR of $k$-th UE is described as
@@
$\underset{\boldsymbol{\theta},\{\mathbf{f}_{l}\}_{l=1}^{L}}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \frac{1}{\sigma_{\rm T}^{2}}\sum_{l=1}^{L}\rho_{l}$ (8a)
@@
where $P$ is the available transmit power of the BS. This problem is evidently non-convex, mainly due to the coupling between the optimization variables and the non-convex unitary modulus constraints. In this section, we will devise an alternating optimization (AO) algorithm, which switches between optimizing the precoding vectors and RIS phase-shifts until convergence.
@@ III-A Precoder optimization
In this subsection, we are going to show the first subproblem of the AO algorithm, that is the problem that retrieves the optimal precoder $\mathbf{F}$ while $\boldsymbol{\Theta}$ is fixed. It can be shown that the sensing SNR per stream, $\rho_{l}$, can be written as $\rho_{l}=\mathbf{f}_{l}^{\mathsf{H}}\boldsymbol{\boldsymbol{\Gamma}}\mathbf{f}_{l}$, where the matrix $\boldsymbol{\Gamma}$ is
$\boldsymbol{\Gamma}$ $=\delta_{s}^{2}\mathbf{H}_{s,{\rm T}}^{\mathsf{H}}\mathbf{H}_{s,{\rm T}}+\delta_{m}\mathbf{H}_{s,{\rm T}}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}$ (9)
$\quad+\delta_{m}^{*}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}^{\mathsf{H}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{H}_{s,{\rm T}}$
$\quad+\delta_{r}^{2}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}^{\mathsf{H}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}$
@@
$\text{P}1:=\underset{\mathbf{W}_{l}\succeq 0,\forall l}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \frac{1}{\sigma_{\text{T}}^{2}}\sum_{l=1}^{L}\text{Tr}\Big{(}\mathbf{W}_{l}\boldsymbol{\Gamma}\Big{)}$ (10a)
@@
where $\tilde{\mathbf{H}}_{k}=\mathbf{h}_{k}\mathbf{h}_{k}^{\mathsf{H}}$. We can apply semi-definite relaxation (SDR) to P1 by removing the rank one constraint. The resulting problem becomes convex and can be solved using commercially available solvers like CVX. During our simulations, we have observed that, for the scenario at hand, the optimal solutions $\{\mathbf{W}_{l}^{\text{opt}}\}_{l=1}^{K}$ are rank one, thus the relaxation is tight and $\{\mathbf{f}_{l}^{\text{opt}}\}_{l=1}^{K}$ can be retrieved through an eigendecomposition. We cannot say the same about $\mathbf{W}_{L}$, thus, we employ the rand C method from [12], generating a set $\mathcal{Z}$ of randomized solutions that shall be denoted by $\mathbf{f}_{L}^{\text{rd}}$. We further define the set of feasible randomized solutions as $\mathcal{S}=\Big{\{}\mathbf{f}_{L}^{\text{rd}}:\mathbf{f}_{L}^{\text{rd}}\in\mathcal{Z},\leavevmode\nobreak\ \|\mathbf{f}_{L}^{\text{rd}}\|^{2}\leq P-\sum_{k=1}^{K}\|\mathbf{f}_{k}^{\text{opt}}\|^{2},\leavevmode\nobreak\ \text{SINR}_{k}(\boldsymbol{\theta},\tilde{\mathbf{F}}^{\text{opt}})\geq\gamma_{k},\leavevmode\nobreak\ k=1,\dots,K\Big{\}}$, where $\tilde{\mathbf{F}}^{\text{opt}}=[\mathbf{f}_{1}^{\text{opt}}\dots\,\mathbf{f}_{K}^{\text{opt}}\ \mathbf{f}_{L}^{\text{rd}}]$. The sensing precoding vector is then obtained as

### v3

# Destructive and Constructive RIS Beamforming in an ISAC Multi-User MIMO Network

@@ Abstract
Integrated sensing and communication (ISAC) has already established itself as a promising solution to the spectrum scarcity problem, even more so when paired with a reconfigurable intelligent surface (RIS), as RISs can shape the propagation environment by adjusting their phase-shift coefficients. Albeit the potential performance gain, a RIS is also a potential security threat to the system. In this paper, we explore both the positive and negative sides of having a RIS in a multi-user multiple-input multiple-output (MIMO) ISAC network. We first develop an alternating optimization algorithm, obtaining the active and passive beamforming vectors that maximize the sensing signal-to-noise ratio (SNR) under minimum signal-to-interference-plus-noise ratio (SINR) constraints for the communication users and finite power budget. We also investigate the destructive potential of the RIS by devising a RIS phase-shift optimization algorithm that minimizes the sensing SNR while preserving the same minimum communication SINR previously guaranteed by the system. We further investigate the impact of the RIS’s individual element failures on the system performance. The simulation results show that the RIS performance-boosting potential is as good as its destructive one and that both of our optimization strategies are hindered by the investigated impairments.
@@ I Introduction
Reconfigurable intelligent surfaces have captured significant interest from both academia and industry for their ability to control challenging radio propagation environments intelligently. By using low-cost passive reflecting elements that induce a controllable phase-shift to incoming waveforms, RISs can effectively steer reflected signals towards desired locations while reducing interference at undesired points [1, 2]. Moreover, RISs have recently emerged as a promising technology in integrated sensing and communications (ISAC) applications, increasing the available spatial degrees of freedom. In [3], the sum rate of the communication user equipments is maximized under the worst-case sensing signal-to-noise ratio (SNR) constraint. On the other hand, [4] defines a sensing signal-to-interference-plus-noise ratio (SINR) by taking into account the presence of interfering objects along the sensing path: this sensing SINR is then maximized in a weighted fashion alongside minimizing the multi-user interference.
Despite their cost-effective deployment and potential to enhance wireless communication and sensing performance, RISs present critical physical layer security (PLS) challenges. In particular, RISs could potentially be exploited as malicious entities that instead compromise communication and sensing performance. [5] gives an overview of the possible ways in which an illegally deployed RIS can hinder system performance, characterizing signal leakage from the legitimate system to the said RIS and the potential damage of interference attacks launched by the former. Adding artificial noise onto the legitimate system’s transmitted waveform proves to be an effective strategy in increasing the system’s PLS, even though the impact of an illegally deployed RIS is still non-negligible.
@@
On the other hand, [6] assesses the impact of an active RIS, i.e., a RIS capable of amplifying the impinging signal, onto an Internet of Things network where the RIS aims at minimizing a single device’s SNR. The authors analyze both the cases of known and unknown malicious RIS identity. A significant challenge is that the attacks are difficult to detect because they are not carried out by generating interfering signals, as in traditional jamming. Instead, they involve configuring the RIS to degrade the system’s performance. In [7], the impact of these silent attacks is demonstrated, further showing how destructive beamforming is susceptible to channel state information (CSI) uncertainties. An important albeit often overlooked practical consideration about RIS is that individual elements, also known as pixels, may fail: [8] gives an overview of the different types of errors that may arise in a practical RIS implementation and characterizes their impact on the RIS radiation pattern. In [9], the authors propose a failure model to specify the amplitude and phase-shift of faulty elements and propose a diagnostic method while the authors of [11] assess the impact of RIS pixel failures onto downlink localization.
In this paper, we consider a multi-user (MU)- multiple-input multiple-output (MIMO) network equipped with an ISAC base station (BS), serving multiple UEs while sensing a target. We first characterize the potential positive impact of a RIS deployment by devising a sensing-centric optimization problem computing active and passive beamforming vectors that maximize the sensing SNR under UEs’ SINR constraints and a finite power budget. Inspired by [10], this problem is solved through alternating optimization. We then investigate the RIS destructive potential when, unbeknownst to the BS, the sensing SNR is minimized while the same minimum SINR is still guaranteed to the UEs, to make this attack harder to detect. We further define a pixel fault model and assess its impact on both optimization algorithms. Numerical simulations show that the RIS is equally capable of boosting and degrading the system performances and how the performances of both algorithms are hindered by the investigated pixel failures.
@@ II System Model
We consider the RIS-assisted MU-MIMO ISAC network shown in Fig. 1, where a BS, equipped with $M$ transmit and $M$ receiving antennas and possessing digital beamforming capabilities, serves $K$ single-antenna UEs in the downlink while sensing the presence of a target. The system is equipped with a RIS made of $N$ reflective elements, whose phase-shifts are described by the vector $\boldsymbol{\theta}=[\theta_{1}\dots\theta_{N}]^{\top}$, where $|\theta_{n}|=1$ for $n=1,\dots,N$. The static path between the BS and UE $k$ is denoted by $\mathbf{h}_{\textrm{s},k}\in\mathbb{C}^{M}$ whereas the static path between the BS and target is denoted by $\mathbf{h}_{\textrm{s,T}}\in\mathbb{C}^{M}$. The reflected paths are defined in a similar fashion, with $\mathbf{h}_{\textrm{r},k}\in\mathbb{C}^{N}$ connecting the RIS and UE $k$ while the target is connected to the RIS via $\mathbf{g}\in\mathbb{C}^{N}$. Lastly, the BS is connected to the RIS through the channel $\mathbf{H}_{t}\in\mathbb{C}^{N\times M}$. The end-to-end channel between the BS and UE $k$ is defined as
$\mathbf{h}_{k}=\mathbf{h}_{\textrm{s},k}+\mathbf{H}_{t}^{\mathsf{H}}\underbrace{\text{diag}(\boldsymbol{\theta})}_{\triangleq\boldsymbol{\Theta}}\mathbf{h}_{\textrm{r},k}.$ (1)
@@
where $\mathbf{f}_{l}\in\mathbb{C}^{M}$ is the precoding vector associated with the $l$-th stream. We assume that the information meant for UE $k$ is encoded into stream $k$, while an additional stream $K+1$ is used for sensing purposes. Hence, the total number of streams is $L=K+1$. On the other hand, $t_{l}$ is the complex-valued unit-power information or sensing symbol and $w_{k}\sim\mathcal{CN}(0,\sigma_{k}^{2})$ is the complex Gaussian receiver noise with variance $\sigma_{k}^{2}$. As for the sensing performance, the two-way sensing channel is
$\mathbf{H}_{\rm T}=c_{r}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\underbrace{\mathbf{g}\mathbf{g}^{\mathsf{H}}}_{\triangleq\mathbf{T}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}+c_{s}\underbrace{\mathbf{h}_{\textrm{s,T}}\mathbf{h}_{\textrm{s,T}}^{\mathsf{H}}}_{\triangleq\mathbf{h}_{\textrm{s,T}}}.$ (3)
The radar cross sections associated with the two sensing paths, which also accounts for their pathloss, are modeled according to the Swerling-I model, that is, $c_{r}\sim\mathcal{CN}(0,\delta_{r}^{2})$ and $c_{s}\sim\mathcal{CN}(0,\delta_{s}^{2})$ and they take a single value throughout the transmission. The monostatic sensing observation can be modeled as
@@
where $\mathbf{F}=[\mathbf{f}_{1}\dots\mathbf{f}_{L}]$ , $\mathbf{t}=[t_{1}\dots t_{L}]^{\top}$, and the entries of $\mathbf{w}_{\rm T}$ are independent and distributed as $\mathcal{CN}(0,\sigma_{\rm T}^{2})$. As for the RIS’s pixel malfunctioning, we adopt the clustered-biased model defined in [8]: the bias may be caused by bit-flip errors at the computing plane of the RIS or by external environmental factors, which can in turn cause a cascade effect onto the neighboring pixel’s shifters, leading to the clustered nature of the failure. We collect the neighboring faulty pixels in the set $\mathcal{Q}$. The affected pixels assume a phase-shift at a fixed distance $\kappa$ from the required one. This is implemented by defining a failure mask $\mathbf{m}=[m_{1}\dots m_{N}]^{\top}$ as [11]
$m_{n}=\begin{cases}e^{j\kappa},&\text{if}\leavevmode\nobreak\ n\in\mathcal{Q},\\ 1,&\text{otherwise}.\end{cases}.$ (5)
@@ III Legitimate Sensing SNR Maximization
The legitimate reason for deploying the RIS is to improve the system performance. Hence, the BS and RIS jointly maximize the sensing SNR while guaranteeing a minimum SINR to the UEs. The choice of optimization function is motivated by the fact that the sensing SNR is expected to have a direct proportionality to various sensing tasks, such as detection and tracking. More specifically,[cheng2023coordinated, cheng2024optimal] show that the detection probability monotonically increases with the sensing SNR. The sensing SNR can be defined as
$\rho=\frac{\sum_{l=1}^{L}\rho_{l}}{\sigma_{\rm T}^{2}}=\frac{\sum_{l=1}^{L}\mathbb{E}\left\{\left\lVert\mathbf{H}_{\rm T}\mathbf{f}_{l}\right\rVert^{2}\right\}}{\sigma_{\rm T}^{2}}$ (6)
Similarly, the SINR of the $k$-th UE is described as
@@
$\underset{\boldsymbol{\theta},\{\mathbf{f}_{l}\}_{l=1}^{L}}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \frac{\sum_{l=1}^{L}\rho_{l}}{\sigma_{\rm T}^{2}}$ (8a)
@@
where $P$ is the available transmit power of the BS. This problem is non-convex, mainly due to the coupling between the optimization variables and the non-convex unitary modulus constraints. In this section, we will devise an alternating optimization (AO) algorithm, which switches between optimizing the precoding vectors and RIS phase-shifts until convergence.
@@ III-A Precoder optimization
In this subsection, we will consider the first subproblem of the AO algorithm, which is the problem that retrieves the optimal precoder $\mathbf{F}$ when $\boldsymbol{\Theta}$ is fixed. It can be shown that the sensing SNR per stream, $\rho_{l}$, can be written as $\rho_{l}=\mathbf{f}_{l}^{\mathsf{H}}\boldsymbol{\boldsymbol{\Gamma}}\mathbf{f}_{l}$, where the matrix $\boldsymbol{\Gamma}$ is
$\boldsymbol{\Gamma}$ $=\delta_{s}^{2}\mathbf{h}_{\textrm{s,T}}^{\mathsf{H}}\mathbf{h}_{\textrm{s,T}}+\delta_{m}\mathbf{h}_{\textrm{s,T}}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}$
$\quad+\delta_{m}^{*}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}^{\mathsf{H}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{h}_{\textrm{s,T}}$
$\quad+\delta_{r}^{2}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}^{\mathsf{H}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{H}_{t}^{\mathsf{H}}\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}$ (9)
@@
$\text{P}1:=\underset{\mathbf{W}_{l}\succeq 0,\forall l}{\mathrm{maximize}}\,\,$ $\leavevmode\nobreak\ \frac{\sum_{l=1}^{L}\text{Tr}\Big{(}\mathbf{W}_{l}\boldsymbol{\Gamma}\Big{)}}{\sigma_{\text{T}}^{2}}$ (10a)
@@
where $\tilde{\mathbf{H}}_{k}=\mathbf{h}_{k}\mathbf{h}_{k}^{\mathsf{H}}$. We can apply semi-definite relaxation (SDR) to P1 by removing the rank one constraint. The resulting problem becomes convex and can be solved using general-purpose solvers like CVX. During our simulations, we have observed that, for the scenario at hand, the optimal solutions $\{\mathbf{W}_{l}^{\text{opt}}\}_{l=1}^{K}$ have rank one. Thus the relaxation is tight and $\{\mathbf{f}_{l}^{\text{opt}}\}_{l=1}^{K}$ can be retrieved through an eigendecomposition. We cannot say the same about $\mathbf{W}_{L}$, thus, we employ the rand C method from [12], generating a set $\mathcal{Z}$ of randomized solutions that shall be denoted by $\mathbf{f}_{L}^{\text{rd}}$. We further define the set of feasible randomized solutions as $\mathcal{S}=\{\mathbf{f}_{L}^{\text{rd}}:\mathbf{f}_{L}^{\text{rd}}\in\mathcal{Z},\leavevmode\nobreak\ \|\mathbf{f}_{L}^{\text{rd}}\|^{2}\leq P-\sum_{k=1}^{K}\|\mathbf{f}_{k}^{\text{opt}}\|^{2},\leavevmode\nobreak\ \text{SINR}_{k}(\boldsymbol{\theta},\tilde{\mathbf{F}}^{\text{opt}})\geq\gamma_{k},\leavevmode\nobreak\ k=1,\dots,K\}$, where $\tilde{\mathbf{F}}^{\text{opt}}=[\mathbf{f}_{1}^{\text{opt}}\dots\,\mathbf{f}_{K}^{\text{opt}}\ \mathbf{f}_{L}^{\text{rd}}]$. The sensing precoding vector is then obtained as